i have a pie chart that i need to complete. However, it is all in percentages and i need to covert those percentages in to numbers to work out the total number and draw the pie chart. laundry is 20%. Toilet flushing is 30%. Kitchen and drinking is 10%. Shower and bath is 35%. Cleaning is 5%. please answer quickly as i need to hand this in tomorrow.
thanks

Answer by Edwin McCravy(20086)

i have a pie chart that i need to complete. However, it is all in percentages and i need to covert those percentages in to numbers to work out the total number and draw the pie chart.
Water usage            percentage      decimal×360°  = central angle 

laundry                   20%             .20×360°   =    72° 
Toilet flushing           30%             .30×360°   =   108°
Kitchen and drinking      10%             .10×360°   =    36°
Shower and bath           35%             .35×360°   =   126°    
Cleaning                   5%             .05×360°   =    18°
------------------------------------------------------------------
                         100%                            360°

Be sure to measure those central angles with a protractor.  



Edwin